# Characterizing Dark Bosons at Chiral Belle

### v1

@@ Abstract
We explore the advantages of a polarized electron beam at Belle II, as proposed for “Chiral Belle”, in the search for invisibly decaying (dark) bosons that weakly couple to the Standard Model. By measuring the polarization dependence of the production cross section of dark bosons in association with a photon, the dark boson’s spin and Lorentz structure of its couplings can potentially be determined. We analyze the mono-photon channel, $e^{+}e^{-}\rightarrow\gamma+\text{invisible}$, in detail, focusing on the production of an on-shell spin-1 boson. We explore this in the context of three separate scenarios for a new dark vector: a dark photon, a mass-mixed “dark $Z$”, and a vector that couples to right-handed electrons, and estimate how well the couplings of such bosons to electrons can be constrained in the event of a positive signal.
@@ I Introduction
Despite its many successes, there is strong reason to believe that the Standard Model (SM) of particle physics is not the last word when it comes to a fundamental understanding of nature. One of the most pressing deficiencies of the SM is its lack of a candidate for the dark matter observed in the Universe. There is overwhelming evidence for dark matter astrophysically and cosmologically [1, *Hamilton:2000du, *Bahcall_2003, *Allen_2011, *Simon_2019], while thus far there has been no conclusive sign of dark matter in Earth-based searches.
Dark matter (DM) so far has only been probed gravitationally, and little is known about its fundamental nature, in particular, its mass and non-gravitational interactions. Dark matter candidates have been proposed over a wide range of masses, from $\sim 10^{-20}~{}\rm eV$ to the Planck scale or even larger if the dark matter is a macroscopic object [6, *Lin:2019uvt, *Lisanti:2016jxe, *Bertone:2004pz, *Jungman:1995df]. Given this ignorance, it is important that we cast a wide net in our search for DM and thoroughly explore the mass scales that we have access to experimentally.
In recent years, DM with a mass between about $10~{}\rm MeV$ and a few $\rm GeV$ has received increased attention [11, *Alexander:2016aln, *Battaglieri:2017aum, *Agrawal:2021dbo, *Gori:2022vri, *deLima:2020lzl]. In this mass range, DM can be difficult to detect in direct detection experiments even if it is not very feebly coupled to the SM [17]. A simple setup to realize such a scenario involves DM coupled to the SM through the vector or scalar portals – that is, DM is coupled to a vector or scalar boson that weakly interacts with SM fields, which we generically term a dark boson. The existence of such a mediator is also motivated to evade the Lee-Weinberg lower bound on sub-GeV DM [18]. If the dark boson is produced at a terrestrial experiment through its small coupling to the SM and its mass is more than twice the DM mass, it is natural that the mediator decays to DM and appears experimentally as missing momentum.
@@
Dark bosons at Belle II have been studied extensively for unpolarized beams [30, 31, 32, 33, 34, 35, 36, 37, 38] . In these studies, it is seen that the expected reach for dark bosons in particular is world-leading over a large mass range, largely because of the incredible integrated luminosity of $50\ \text{ab}^{-1}$ that Belle II is expected to reach by the end of its running. Recently, it was pointed out that it is possible to distinguish the spin of a dark boson with Chiral Belle [37].
@@
In this work, we focus on light dark vectors ($<10$ GeV) which decay on-shell to invisible states. We work out the polarization dependence of the production of an arbitrary dark boson, $X$, in the mono-photon channel, $e^{+}e^{-}\rightarrow\gamma X$. While a dark (pseudo-)scalar has no polarization dependence, as shown in Appendix A, the vector has a non-trivial dependence on beam polarization in the presence of both vector and axial-vector couplings. This analysis is further extended for off-shell decays and heavier masses in [40].
We highlight that the choice of beam polarization dictates which type of new physics the experiment is most sensitive to. Our study considers electron beam polarizations of ±$\pm 70\%$, in line with current Chiral Belle proposals [22]. We further show how the ability to separate the data according to $e^{-}$ polarization opens up the possibility of determining the Lorentz structure of such a dark vector in the event of a positive signal. Our study helps identify the major detector-related limitations to searches of this type, notably gaps in the detector coverage and photon reconstruction inefficiencies.
@@ II.1 Kinetically mixed dark photon
A particularly simple way to couple a vector boson to the SM is to kinetically mix it with the $U(1)$ hypercharge gauge boson [41, *Galison:1983pa, *Holdom:1985ag]
@@ II.2 Mass mixed dark $Z$
Another possibility to weakly couple a new massive spin-1 boson to the SM is to allow for mass mixing with the SM $Z$ boson, a scenario that leads to a “dark $Z$” [45]. The relevant term in the Lagrangian is
@@
where $\epsilon_{Z}$ is a small mixing parameter that can be generated in, e.g., an extended Higgs sector. In the two-Higgs doublet-model of [45], $\epsilon_{Z}\simeq\delta\times m_{V}/m_{Z}$ with $\delta$ a parameter of the theory. Atomic parity violation (APV) measurements limit $\delta\lesssim 10^{-2}$ [46, *Porsev:2010de]. In this scenario, the dark $Z$ boson couplings to SM fermions are $Z$-like, and, in particular for the electron, are given by
$g_{V}$ $=\frac{e}{\sin 2\theta_{W}}\left(-\frac{1}{2}+2\sin^{2}\theta_{W}\right)\epsilon_{Z}\simeq-0.05\,\epsilon_{Z}e,$ (5)
$g_{A}$ $=-\frac{e}{\sin 2\theta_{W}}\left(-\frac{1}{2}\right)\epsilon_{Z}\simeq 0.6\,\epsilon_{Z}e\simeq-13g_{V},$
@@ II.3 Right-handed vector
Vectors that couple to right-handed SM fermions have been considered to explain a number of experimental anomalies [49, 50]. Generally, such vectors require additional fermions to render the model theoretically self-consistent, potentially leading to other signals [51, *Batra:2005rh, *Dobrescu:2014fca]. Moreover, since these vectors couple to non-conserved currents at low scales, the emission of longitudinally polarized vectors is enhanced and can lead to very stringent constraints (although there is some model dependence on the UV content of the theory) [54, 55]. In Ref. [55], it was shown that searches for $B^{+}\to K^{+}+{\rm inv.}$ restrict
$\left|g_{V}+g_{A}\right|$ $\lesssim 6\times 10^{-6}\left(\frac{m_{V}}{\rm GeV}\right),$ (6)
for $m_{V}<m_{B}-m_{K}$ in the case of an invisibly decaying vector coupled universally to right-handed SM fermions [56]. Similar to the dark $Z$ scenario, APV in such a model limits
$\left|g_{V}+g_{A}\right|$ $\lesssim 1\times 10^{-4}\left(\frac{m_{V}}{\rm GeV}\right),$ (7)
@@
It is worth noting that gauging left-handed charged leptons is significantly more challenging. It generally leads to enhanced interactions of neutrinos with matter, which are heavily constrained experimentally [49].
@@ II.4 Generic dark vector
$g_{A}$ $\simeq 0.6\,\epsilon_{Z}e\simeq-0.6\times\frac{\epsilon_{Z}}{\epsilon}\times g_{V},$
@@ III Dark Vectors at (Chiral) Belle II
$=1.25~{}{\rm nb}\times\frac{2}{\sqrt{s}}\frac{{E_{V}^{\ast}}^{2}+{E_{\gamma}^{\ast}}^{2}\cos^{2}\theta_{\gamma}^{\ast}}{{E_{\gamma}^{\ast}}\sin^{2}\theta_{\gamma}^{\ast}},$
@@
where $P_{L,R}$ is the polarization defined in Eq. (13) for the nominally left- or right-polarized $e^{-}$ beam and we have used superscripts to specify the couplings of $V$ to the electron and $f$, writing the electric charge of $f$ as $Q_{f}$ where $Q_{e}=-1$.
@@ IV.1 Analysis
We perform our analysis for both an unpolarized and a ±$\pm 70\%$ polarized electron beam. In both the unpolarized and polarized cases, the signal for our analysis is a single hard photon, as we assume that the dark vector decays invisibly. To avoid inefficiencies at lower photon energies, we consider only photons with center of mass energy $E_{\gamma}^{*}>1.8$ GeV. The dominant backgrounds are electromagnetic and independent of the electron beam polarization at leading order. They are
@@
This consists of the same final state as the signal: a single photon recoiling against missing momentum. The only difference between this and the signal process is the virtuality carried by the $\nu_{l}\bar{\nu}_{l}$ pair. This background thus features a range of center-of-mass energies carried by the photon, in contrast to a monochromatic photon in as the signal case (in the center-of-mass frame). This irreducible background is an interesting electroweak physics target in its own right, and is explored further in [40].
@@
We generate the background events for both unpolarized and polarized $e^{-}$ beam at leading order in QED using MadGraph5_aMC@NLO [65] interfaced with Pythia8.2 [66] for showering and hadronization. Because QED processes have large cross sections at this energy, we generate 20 million Monte Carlo events to ensure that we do not miss any relevant kinematic features. In Fig. 1, we show the distribution of the backgrounds with $20~{}{\rm fb}^{-1}$ of data in the $E_{\gamma}^{\ast}$ vs $\theta^{\text{lab}}_{\gamma}$ plane.
Figure 1: Distribution of the mono-photon SM background for 20 fb-1 of data as a function of the center-of-mass energy of the measured single photon $E_{\gamma}^{\ast}$, and the photon lab-frame angle $\theta_{\gamma}^{\text{lab}}$ relative to the electron beam, with the ECL geometry information included. The arch shapes are a consequence of the endcap gaps (which are visible in the histogram) that give a significantly higher probability of missing a particle there and still measuring a photon in the ECL. The concentration of events at high energies comes from $\gamma\gamma$ events where one photon is missed inside the main barrel of the ECL.
@@
Figure 2: Left: Estimated Belle II 95% CL upper limits on $\epsilon_{\rm eff}$ as a function of $m_{V}$ given the analysis described in the text, for both unpolarized and polarized beams with a luminosity 20 fb-1 (black curve), a polarized electron beam with $P=0.7$ and a luminosity of 20 ab-1 (blue curve), and a polarized electron beam with $P=-0.7$ and a luminosity of 20 ab-1 (orange curve). The bound from polarized beams should be interpreted as being on $\epsilon_{\rm eff}$ as defined in Eq. (14). Existing limits for invisible dark photon searches from E787/E949 [30, 67, *BNL-E949:2009dza, *Davoudiasl:2014kua], BaBar [60], NA62 [70], BESIII [71], NA64 [72], and the electron anomalous magnetic moment $a_{e}$ [73] (assuming a vector coupling and the measurement of $\alpha$ in Ref. [74]—taking an axial-vector coupling and the larger $\alpha$ in Ref. [75] does not change the bound substantially). The gray-dashed curves are thermal DM targets with $m_{V}=3m_{\rm DM}$ and $\alpha_{D}=0.5$ for scalar, Majorana, and pseudo-Dirac DM candidates [76]. Right: The 95% CL limits in the $g_{V}$ vs $g_{A}$ plane at Chiral Belle, assuming 20 ab-1 of $70\%$ left-handed (orange) or $70\%$ right-handed (blue) polarized electron beams (regions outside the ellipses are excluded). We also show the Belle II reach for unpolarized beams with 40 ab-1 of data (black). The dashed green lines indicate the relationship between $g_{A}$ and $g_{V}$ for a dark photon, a dark $Z$ boson, and a right-handed vector as labeled.
@@
We first consider an ideal detector with $20~{}\rm fb^{-1}$ of unpolarized data and a $100$ MeV dark vector. The signal region is localized at $E_{\gamma}^{*}=5.29~{}\rm GeV$, and we construct an energy bin from $4.89~{}\rm GeV$ to $5.29~{}\rm GeV$. The optimal cut that maximizes the $95\%$ CL exclusion sensitivity is $\theta_{\gamma}^{\rm lab}=[39^{\circ},112^{\circ}]$ and inside this region we expect no background events ($b=0.13$). This translates into an estimated 95% CL limit $\epsilon_{\rm eff}<2.04\times 10^{-4}$. Once we include smearing based on the detector energy resolution, events from lower energies bleed into the high energy region, and the optimal signal region shrinks to $\theta_{\gamma}^{\rm lab}=[58^{\circ},102^{\circ}]$. We obtain a weakened limit of $\epsilon_{\rm eff}<2.61\times 10^{-4}$. Finally, including a photon detection inefficiency of $10^{-6}$ in the main barrel of the ECL does not affect the optimal signal region but weakens the limit even further to $\epsilon_{\rm eff}<4.90\times 10^{-4}$.
@@
The reach of Belle II with $20~{}\rm fb^{-1}$ of unpolarized or polarized data is depicted by the black curve. Note that the reach does not depend on the polarization since the neutrino background is not yet significant. We see that Belle II covers new parameter space compared to BaBar for $m_{V}\lesssim 3$ GeV. The reach for $P=0.7$ with $20~{}\rm ab^{-1}$ of data is shown in the left plot of Fig. 2 by the blue curve, while the reach for $P=-0.7$ is shown by the dashed orange curve. In both cases, dark vector couplings $\lesssim 10^{-4}$ will be probed over a broad range of dark vector masses. We see that the reach for either polarization is essentially equal. This is because the dominant background is not sensitive to polarization and comes from the photon inefficiency. The subdominant neutrino background does depend on polarization, and using polarization information aids in measuring this background as well as in searches for off-shell dark vectors as we explore in Ref. [40].
Figure 3: Example of the discriminatory power from $e^{-}$ beam polarization for the three different models we consider. In these plots, we assume a future $5\sigma$ discovery with with $40~{}{\rm ab}^{-1}$ of unpolarized data for $m_{V}=0.1~{}\rm GeV$ which carves out the gray circular $1\sigma$ preferred coupling regions. The $20~{}{\rm ab}^{-1}$ of $P=0.7$ and $P=-0.7$ (orange) data in this sample then require at $1\sigma$ that the couplings are in the blue and orange regions respectively, (partially) breaking the degeneracy in the $g_{V}$ and $g_{A}$ plane. The dashed green lines are the coupling relationships for each scenario. The red shaded regions show the couplings preferred at the $1\sigma$ level from measurements of $A_{LR}$ defined in Eq. (18), which is less subject to systematic errors. Note that there is a degeneracy in the $A^{\prime}$ and $Z_{d}$ cases between (nearly) purely vector or axial-vector couplings so that the cross sections for left- or right-polarized $e^{-}$ beams are nearly the same. In the $Z_{R}$ case, the chiral nature of the couplings breaks this degeneracy.

### v2

@@ Abstract
We explore the advantages of a polarized electron beam at Belle II, as proposed for “Chiral Belle”, in the search for invisibly decaying (dark) bosons that weakly couple to the Standard Model. By measuring the polarization dependence of the production cross section of dark bosons in association with a photon, the dark boson’s spin and Lorentz structure of its couplings can potentially be determined. We analyze the mono-photon channel, $e^{+}e^{-}\rightarrow\gamma+\text{invisible}$, in detail, focusing on the production of an on-shell spin-1 boson. We explore this in the context of three separate scenarios for a new dark vector: a dark photon, a mass-mixed “dark ZZ”, and a vector that couples to right-handed electrons, and estimate how well the couplings of such bosons to electrons can be constrained in the event of a positive signal.
@@ I Introduction
In defiance of its many successes, there is strong reason to believe that the Standard Model (SM) of particle physics is not the last word when it comes to a fundamental understanding of nature. One of the most pressing deficiencies of the SM is its lack of a candidate for the dark matter observed in the Universe. There is overwhelming evidence for dark matter astrophysically and cosmologically [59, *Hamilton:2000du, *Bahcall_2003, *Allen_2011, *Simon_2019], while thus far there has been no conclusive sign of dark matter in Earth-based searches.
Dark matter (DM) so far has only been probed gravitationally, and little is known about its fundamental nature, in particular, its mass and non-gravitational interactions. Dark matter candidates have been proposed over a wide range of masses, from $\sim 10^{-20}~\rm eV$ to the Planck scale or even larger if the dark matter is a macroscopic object [23, *Lin:2019uvt, *Lisanti:2016jxe, *Bertone:2004pz, *Jungman:1995df]. Given this ignorance, it is important that we cast a wide net in our search for DM and thoroughly explore the mass scales that we have access to experimentally.
In recent years, DM with a mass between about $10~\rm MeV$ and a few $\rm GeV$ has received increased attention [34, *Alexander:2016aln, *Battaglieri:2017aum, *Agrawal:2021dbo, *Gori:2022vri, *deLima:2020lzl]. In this mass range, DM can be difficult to detect in direct detection experiments even if it is not very feebly coupled to the SM [60]. A simple setup to realize such a scenario involves DM coupled to the SM through the vector or scalar portals – that is, DM is coupled to a vector or scalar boson that weakly interacts with SM fields, which we generically term a dark boson. The existence of such a mediator is also motivated to evade the Lee-Weinberg lower bound on sub-GeV DM [45]. If the dark boson is produced at a terrestrial experiment through its small coupling to the SM and its mass is more than twice the DM mass, it is natural that the mediator decays to DM and appears experimentally as missing momentum.
@@
[…]
@@
In this work, we focus on light dark vectors (<10<10 GeV) which decay on-shell to invisible states. We work out the polarization dependence of the production of an arbitrary dark boson, $X$, in the mono-photon channel, $e^{+}e^{-}\rightarrow\gamma X$. While a dark (pseudo-)scalar has no polarization dependence, as shown in Appendix A, the vector has a non-trivial dependence on beam polarization in the presence of both vector and axial-vector couplings. This analysis is further extended for off-shell decays and heavier masses in [30].
We highlight that the choice of beam polarization dictates which type of new physics the experiment is most sensitive to. Our study considers electron beam polarizations of $\pm 70\%$, in line with current Chiral Belle proposals [14]. We further show how the ability to separate the data according to $e^{-}$ polarization opens up the possibility of determining the Lorentz structure of such a dark vector in the event of a positive signal. Our study helps identify the major detector-related limitations to searches of this type, notably gaps in the detector coverage and photon reconstruction inefficiencies.
@@ II.1 Kinetically mixed dark photon
A particularly simple way to couple a vector boson to the SM is to kinetically mix it with the $U(1)$ hypercharge gauge boson [53, *Galison:1983pa, *Holdom:1985ag]
@@ II.2 Mass mixed dark $Z$
Another possibility to weakly couple a new massive spin-1 boson to the SM is to allow for mass mixing with the SM $Z$ boson, a scenario that leads to a “dark ZZ” [29]. The relevant term in the Lagrangian is
@@
where $\epsilon_{Z}$ is a small mixing parameter that can be generated in, e.g., an extended Higgs sector. In the two-Higgs doublet-model of [29], $\epsilon_{Z}\simeq\delta\times m_{V}/m_{Z}$ with $\delta$ a parameter of the theory. Atomic parity violation (APV) measurements limit $\delta\lesssim 10^{-2}$ [49, *Porsev:2010de]. In this scenario, the dark $Z$ boson couplings to SM fermions are $Z$-like, and, in particular for the electron, are given by
[…]
@@ II.3 Right-handed vector
Vectors that couple to right-handed SM fermions have been considered to explain a number of experimental anomalies [16, 44]. Generally, such vectors require additional fermions to render the model theoretically self-consistent, potentially leading to other signals [55, *Batra:2005rh, *Dobrescu:2014fca]. Moreover, since these vectors couple to non-conserved currents at low scales, the emission of longitudinally polarized vectors is enhanced and can lead to very stringent constraints [32, 31] although there is strong model dependence on the anomaly-canceling UV content of the theory [37]. In Ref. [31], it was shown that a limit on the branching ratio for $B^{+}\to K^{+}+{\rm inv.}$ of $2.5\times 10^{-5}$ restricts
[…]
for $m_{V}<m_{B}-m_{K}$ in the case of an invisibly decaying vector coupled universally to right-handed SM fermions [43]. There is now a measurement from Belle II of the branching fraction for this mode of $(2.3\pm 0.7)\times 10^{-5}$ [7] which differs at a more than $2\sigma$ from the SM prediction; for a discussion of potential explanations of this see, e.g., [11, *McKeen:2023uzo, *Fridell:2023ssf, *He:2024iju, *Buras:2024ewl, *Altmannshofer:2024kxb, *Bhattacharya:2024clv, *Calibbi:2025rpx]. For $m_{V}<m_{K}-m_{\pi}\simeq 350~{\rm MeV}$, the limit from kaon decays on the size of the coupling is stronger by about an order of magnitude.
Similar to the dark $Z$ scenario, APV in such a model limits
[…]
@@
It is worth noting that gauging left-handed charged leptons is significantly more challenging. It generally leads to enhanced interactions of neutrinos with matter, which are heavily constrained experimentally [16]. Additionally, our study below will focus on the coupling of the new boson to electrons and not make use of the model-dependent limits above that rely on additional couplings to quarks.
@@ II.4 Generic dark vector
[…]
@@ III Dark Vectors at (Chiral) Belle II
[…]
@@
where $P_{L,R}$ is the polarization defined in Eq. (13) for the nominally left- or right-polarized $e^{-}$ beam and we have used superscripts to specify the couplings of $V$ to the electron and $f$, writing the electric charge of $f$ as $Q_{f}$ where $Q_{e}=-1$. (When the couplings and mass are rescaled appropriately, this matches the well-known SM value involving $Z$ exchange in, e.g., Ref. [56].)
@@
Figure 1: Truth-level (left) and smeared (right) distribution of the mono-photon SM background for 20 fb-1 of data as a function of the center-of-mass energy of the measured single photon $E_{\gamma}^{\ast}$, and the photon lab-frame angle $\theta_{\gamma}^{\text{lab}}$ relative to the electron beam, with the ECL geometry information included. The arch shapes are a consequence of the endcap gaps (which are visible in the histogram) that give a significantly higher probability of missing a particle there and still measuring a photon in the ECL. The concentration of events at high energies comes from $\gamma\gamma$ events where one photon is missed inside the main barrel of the ECL.
@@ IV.1 Analysis
We perform our analysis for both an unpolarized and a $\pm 70\%$ polarized electron beam. In both the unpolarized and polarized cases, the signal for our analysis is a single hard photon, as we assume that the dark vector decays invisibly. To avoid inefficiencies at lower photon energies, we consider only photons with center of mass energy $E_{\gamma}^{*}>1.8$ GeV. The dominant backgrounds are electromagnetic and independent of the electron beam polarization at leading order. They are
@@
This consists of the same final state as the signal: a single photon recoiling against missing momentum. The only difference between this and the signal process is the virtuality carried by the $\nu_{l}\bar{\nu}_{l}$ pair. This background thus features a range of center-of-mass energies carried by the photon, in contrast to a monochromatic photon as in the signal case (in the center-of-mass frame). This irreducible background is an interesting electroweak physics target in its own right, and is explored further in [30].
@@
Figure 2: Left: Estimated Belle II 95% CL upper limits on $\epsilon_{\rm eff}$ as a function of $m_{V}$ given the analysis described in the text, for both unpolarized and polarized beams with a luminosity 20 fb-1 (black curve), a polarized electron beam with $P=0.7$ and a luminosity of 20 ab-1 (blue curve), and a polarized electron beam with P=0.7P=-0.7 and a luminosity of 20 ab-1 (orange curve). The bound from polarized beams should be interpreted as being on $\epsilon_{\rm eff}$ as defined in Eq. (14). Existing limits for invisible dark photon searches from E787/E949 [33, 8, *BNL-E949:2009dza, *Davoudiasl:2014kua], BaBar [47], NA62 [25], BESIII [3], NA64 [13], and the electron anomalous magnetic moment $a_{e}$ [36] (assuming a vector coupling and the measurement of $\alpha$ in Ref. [51]—taking an axial-vector coupling and the larger $\alpha$ in Ref. [54] does not change the bound substantially). The gray-dashed curves are thermal DM targets with $m_{V}=3m_{\rm DM}$ and $\alpha_{D}=0.5$ for scalar, Majorana, and pseudo-Dirac DM candidates [20]. Right: The 95% CL limits in the $g_{V}$ vs $g_{A}$ plane at Chiral Belle for $m_{V}=100$ MeV, assuming 20 ab-1 of $70\%$ left-handed (orange) or $70\%$ right-handed (blue) polarized electron beams (regions outside the ellipses are excluded). We also show the Belle II reach for unpolarized beams with 40 ab-1 of data (black). The dashed green lines indicate the relationship between $g_{A}$ and $g_{V}$ for a dark photon, a dark $Z$ boson, and a right-handed vector as labeled.
@@
We generate the background events for both unpolarized and polarized $e^{-}$ beam at leading order in QED using MadGraph5_aMC@NLO [12] interfaced with Pythia8.2 [61] for showering and hadronization. Because QED processes have large cross sections at this energy, we generate 90 million Monte Carlo events to ensure that we do not miss any relevant kinematic features. In Fig. 1, we show the distribution of the backgrounds with $20~{\rm fb}^{-1}$ of data in the $E_{\gamma}^{\ast}$ vs $\theta^{\text{lab}}_{\gamma}$ plane.
@@
Figure 3: Example of the discriminatory power from $e^{-}$ beam polarization for the three different models we consider. In these plots, we assume a future $5\sigma$ discovery with $40~{\rm ab}^{-1}$ of unpolarized data for $m_{V}=3.67~\rm GeV$, which carves out the gray circular $1\sigma$ preferred coupling regions. The $20~{\rm ab}^{-1}$ of $P=0.7$ and P=0.7P=-0.7 (orange) data in this sample then require at $1\sigma$ that the couplings are in the blue and orange regions respectively, (partially) breaking the degeneracy in the $g_{V}$ and $g_{A}$ plane. The dashed green lines are the coupling relationships for each scenario. The red shaded regions show the couplings preferred at the $1\sigma$ level from measurements of $A_{LR}$ defined in Eq. (18), which is less subject to systematic errors. Note that there is a degeneracy in the $A^{\prime}$ and $Z_{d}$ cases between (nearly) purely vector or axial-vector couplings so that the cross sections for left- or right-polarized $e^{-}$ beams are nearly the same. In the $Z_{R}$ case, the chiral nature of the couplings breaks this degeneracy.
@@
We first consider an ideal detector with $20~\rm fb^{-1}$ of unpolarized data and a $100$ MeV dark vector. The signal region is localized at $E_{\gamma}^{*}=5.29~\rm GeV$, and we construct an energy bin from $4.89~\rm GeV$ to $5.29~\rm GeV$. The optimal cut that maximizes the $95\%$ CL exclusion sensitivity is $\theta_{\gamma}^{\rm lab}=[38^{\circ},115^{\circ}]$ and inside this region we expect no background events. This translates into an estimated 95% CL limit $\epsilon_{\rm eff}<1.68\times 10^{-4}$. Once we include smearing based on the detector energy resolution, events from lower energies bleed into the high energy region. We obtain a slightly weakened limit of $\epsilon_{\rm eff}<1.74\times 10^{-4}$. Finally, including a photon detection inefficiency of $10^{-6}$ in the main barrel of the ECL does not affect the optimal signal region but significantly weakens the limit to $\epsilon_{\rm eff}<4.20\times 10^{-4}$.
@@
The reach of Belle II with $20~\rm fb^{-1}$ of unpolarized or polarized data is depicted by the black curve. Note that the reach does not depend on the polarization since the neutrino background is not significant with this luminosity. We see that Belle II covers new parameter space compared to BaBar for $m_{V}\lesssim 3$ GeV. The reach for $P=0.7$ with $20~\rm ab^{-1}$ of data is shown in the left plot of Fig. 2 by the blue curve, while the reach for P=0.7P=-0.7 is shown by the dashed orange curve. In both cases, dark vector couplings $\lesssim 10^{-4}$ will be probed over a broad range of dark vector masses. We see that, for most of the mass range, the reach for either polarization is essentially equal. This is because the dominant background at corresponding photon energies is due to QED and thus not sensitive to polarization. For $3~{\rm GeV}\lesssim m_{V}\lesssim 4~{\rm GeV}$, the subdominant neutrino background, which does depend on polarization, becomes important and leads to differences in the reach for each polarization. The usefulness of polarization information in actually measuring this SM process is further explored in Ref. [30].